What is the equation for the circle with endpoints of a diameter at (6,6) and (10,12)?
Answer by Greenfinch(383)

Center of circle is at (10+ 6)/2 = 8 and (12 + 6)/2 = 9
The distance from one end of the diameter (6,6) to the centre (8,9)
is sqrt 2^2 + 3^2 = sqrt 13
The equation of the circle centrte 8,9 radius sqrt 13 is thus
(x - 8)^2 +(y-9)^2 = 13
x^2 -16x +y^2 - 18y + 81 + 64 = 13
x^2 - 16x + y^2 - 18y + 132 = 0